OpenStudy (anonymous):

How can x2 + 3x + 1 = 2x2 + 2x + 3 be set up as a system of equations?

OpenStudy (anonymous):

first subtract (x2 + 3x + 1) -(2x2 + 2x + 3)=0 which equals to -x^2+x-2 = 0

OpenStudy (anonymous):

then use Quadratic Formula to -x^2+x-2 to solve for x

OpenStudy (anonymous):

let me know if u need more assistance

OpenStudy (anonymous):

how do i plug that into the quadratic formula?

OpenStudy (anonymous):

ax^2 + bx + c.. the formula is x = [ -b ± sqrt(b^2 - 4ac) ] / 2a

OpenStudy (anonymous):

so y plug 1 to b, -1 to a, -2 to c.. then solve it

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

welcome :)